OpenStudy (anonymous):

12/7 = -6y I have to find the value of y Can someone explain?

OpenStudy (anonymous):

\[-6y = \frac{12}{7}\]

OpenStudy (anonymous):

Can you explain?

OpenStudy (anonymous):

Wait, I have not done anything yet.

OpenStudy (anonymous):

See, can you see we want to get \(y\), but \(y\) is not alone, \(-6\) os our enemy, it is not leaving \(y\) for us, getting?

OpenStudy (anonymous):

mhm.. but Im still not exactly sure how to solve..

OpenStudy (anonymous):

For that I am here, just tell me what I am saying to you, are you getting?

OpenStudy (anonymous):

so far, yes

OpenStudy (anonymous):

So, we need to get rid of \(-6\).. See, when we divide by -6, on both the sides, then what will you get?

OpenStudy (anonymous):

I show you: \[\frac{-6y}{-6} = \frac{\frac{12}{7}}{-6}\]

OpenStudy (anonymous):

I have divided by \(-6\) both the sides.

OpenStudy (anonymous):

Okay up to here?

OpenStudy (anonymous):

Mhm

OpenStudy (anonymous):

So, on left hand side, \(-6\) and \(-6\) will get cancelled, giving you your \(y\).

OpenStudy (anonymous):

\[\frac{\cancel{-6}y}{\cancel{-6}} = \frac{\frac{12}{7}}{-6} \implies y= \frac{\frac{12}{7}}{-6}\]

OpenStudy (anonymous):

The right hand side can be simplified more.

OpenStudy (anonymous):

\[y = \frac{\frac{12}{7}}{-6} \implies y = \frac{12}{7 \times (-6)} \implies y = \frac{\cancel{12}^2}{7 \times (- \cancel{6}^1)} \implies \color{green}{y = \frac{-2}{7}}\]